OpenStudy (anonymous):

Select the graph of y = – |x + 2| + 3

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

Since the absolute value of a number is always positive, When we graph the equation y = – |x + 2| + 3 the shape of our graph looks quite different from the graph of y = – (x + 2) + 3. (see the attached graph below for more detail) The graph of every linear absolute value always has an inverted "V" or a sideways "V" shape. This "V" is the result of the negative absolute values becoming positive. The global maximum point is when -|x+2| is 0, that is when x=-2 (A). When y = – |x + 2| + 3 = 0, we can find the two points intersecting the x-axis (B, C).